### v1

# Working Backwards: Learning to Place by Picking

Abstract
We present Learning to Place by Picking (LPP), a method capable of autonomously collecting demonstrations for a family of placing tasks in which objects must be manipulated to specific locations. With LPP, we approach the learning of robotic object placement policies by reversing the grasping process and exploiting the inherent symmetry of the pick and place problems. Specifically, we obtain placing demonstrations from a set of grasp sequences of objects that are initially located at their target placement locations. Our system is capable of collecting hundreds of demonstrations without human intervention by using a combination of tactile sensing and compliant control for grasps. We train a policy directly from visual observations through behaviour cloning, using the autonomously-collected demonstrations. By doing so, the policy can generalize to object placement scenarios outside of the training environment without privileged information (e.g., placing a plate picked up from a table and not at the original placement location). We validate our approach on home robotic scenarios that include dishwasher loading and table setting. Our approach yields robotic placing policies that outperform policies trained with kinesthetic teaching, both in terms of performance and data efficiency, while requiring no human supervision.
I Introduction
Object placing is an essential everyday task that autonomous robot manipulators will need to perform reliably. Many household chores, such as setting the table and loading a dishwasher, for example, require careful placing. Robotic placement involves contact-rich interactions and typically necessitates sensory feedback to determine the state of the in-hand object and the environment. Designing object placement policies is non-trivial and can involve solving a number of complicated subtasks, including object tracking, scene understanding, motion planning, and control.
Imitation learning (IL) provides an appealing, simple, and practical alternative by learning end-to-end control policies from expert demonstrations [1, 2]. The real world is complex and demonstration data should ideally include diverse examples to account for varying factors such as background, lighting conditions, and object types. Some recently-developed robotic systems [3, 4] have demonstrated the use of IL with a large amount of demonstration data to effectively solve a variety of tasks. A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect expert demonstrations.
We propose a novel approach named Learning to Place by Picking (LPP) that facilitates and automates the collection of expert demonstrations for a large subset of placing problems by taking advantage of powerful grasp planners [5]. LABEL:fig:title provides a visual summary of our approach. LPP is a simple self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place problems. Given an environment with objects initially positioned at their target place locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory. In this way, we are able to generate place demonstration data in a self-supervised manner—the picking phase provides the training supervision for the place task. Finally, we learn a placing policy based only on visual observations using the demonstrations collected in the training environment. Crucially, this allows our reactive policy to be used for object placement outside of the training environment and in scenarios where the objects are not initially placed at their target locations (i.e., without privileged information). Our main contributions are:
  1. 1.
    LPP, a self-supervised data collection method for 6-DOF robotic object placements,
  2. 2.
    a flexible grasp planning pipeline that enables a user to specify what to grasp in a scene,
  3. 3.
    a tactile-driven regrasping and compliant control strategy to achieve uninterrupted pick and place data collection without human intervention, and
  4. 4.
    real-world experimental validation of LPP by using the collected demonstrations to train a vision-based policy capable of placing multiple plates in a dishrack and of setting a table.
II Related Work
In this section, we review previous approaches to object placing with manipulators and other work that generally leverage the idea of “working backwards”.
II-A Robotic Manipulator Placing
Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [6, 7, 8, 9], while placing is relatively less explored.
Previous approaches have explored the use of motion planning combined with place-specific objectives [10, 11]. For example, in [10], the authors demonstrate a hierarchical sampling-based motion planner that first finds poses which satisfy constraints (e.g., reachability) and then performs a local optimization for user-specified objectives (e.g., clearance). The place problem has also been studied from the perspective of geometric modelling [12], where the planar surfaces of a discretized object model are matched to planar patches in the environment. Importantly, 3D graphic models of the objects being placed as well as the environment are assumed to be available ahead of time. Other works have looked into combining learning with motion planning to operate directly on point clouds, thereby relaxing the assumption of known object and environment models [13, 14].
Closer to our approach, other works have trained end-to-end place policies using reinforcement learning (RL) [15] and imitation learning (IL) [16, 17]. In [16], the authors demonstrate an inverse reinforcement learning (IRL) technique capable of simultaneously learning a cost function and a control policy from human kinesthetic demonstrations of dish placing. The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in combination with any existing methods that require expert demonstrations.
II-B Working Backwards
The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [18, 19] and in learning visual predictive models for control [20, 21, 22]. Our self-supervised data collection method, which extends the concept of working backwards to the problem of robotic placement, is most similar in spirit to [23, 24, 25]. These methods consider objects that are assumed to be in their desired place configuration during data collection. The manipulator can then pick an object and move back to the initial goal location to generate data. However, we focus on placement as opposed to insertion as our task. Insertion policies generally only reason about local interactions, while placement involves reasoning over a larger scene with the potential for multiple objects and place poses to be involved in the task. Furthermore, unlike prior work [23, 24, 25], our system does not require a human to guide the manipulator to the object and instead leverages tools from grasp planning. Our approach also resembles [26], where time-reversed trajectories are used to generate assembly data from disassembly data in a self-supervised manner. However, we focus on generating demonstration data for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement. In contrast, the self-supervised data collection procedure introduced in [26] uses a 2D open loop pick-and-place action primitive.
III Learning to Place by Picking
We first present an overview of our self-supervised data collection method, LPP, in Section III-A and describe a method for noise augmentation during data collection in Section III-B. We then provide details on the procedure used to train our IL policy in Section III-C.
III-A Self-supervised Data Collection
[…]
Figure 1: The four steps involved in LPP, our autonomous demonstration data collection process for placing. We (a) generate grasps with an off-the-shelf grasp planner [5], (b) compliantly grasp the object to apply minimal forces to the environment while ensuring a stable grasp with tactile sensing, (c) retrieve the object with rotational compliance while storing the trajectory, and (d) generate placement demonstration data by rolling out the reversed grasp trajectories while storing the observations and actions.
LPP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving and (4) placing by reversing; these steps are visualized in Figure 1. During training, the objects of interest are initially in their goal positions.
III-A1 Grasp Planning
[…]
Figure 2: The steps involved in our language-driven grasp planning pipeline. We use (a) Grounding-Dino [27] for object bounding box detection based on user provided texts and (b) Segment Anything [28] on the cropped images to produce object specific masks. We then use (c) Contact-graspnet [5] for grasp generation on only the segmented objects. Top: We specify “green plate” and “blue plate” as the targets for data collection. Bottom: We specify “bowl” and “mug” as the targets for data collection.
At the start of the data collection process, the manipulator moves to a pre-determined pose where the camera has a clear view of the objects of interest in the environment, as shown in Figure 1. We use the grasp planner Contact-GraspNet [5] to generate $L$ candidate grasp poses $\mathbf{T}_{g}=\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in SE(3)$ for all $i\in\{0,...,L\}$. The user then provides text-based descriptions of the objects of interest (e.g., “plates”, “green plate”, “cup” or, more generally, “objects”). Given the user-provided texts, we use Grounding-Dino [27] to find the object bounding boxes. Finally, we run Segment Anything [28] on each of the previously generated image crops from the bounding boxes to generate per-object masks, which Contact-GraspNet can then use to filter grasps that are not of objects of interest: $\bar{\mathbf{T}}_{g}=\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$ and $\bar{\mathbf{T}}_{g}\subseteq\mathbf{T}_{g}$. We visualize an example of this process in Figure 2 for two different scenes. After pruning, we randomly select a single grasp pose, $\mathbf{T}_{g,k}$, from $\bar{\mathbf{T}}_{g}$. To approach the object of interest without collisions, we include a pregrasp pose, $\mathbf{T}^{\text{pre}}_{g,k}$, that is defined with a translational offset above the grasp pose $\mathbf{T}_{g,k}$.
III-A2 Grasping
Given the computed pregrasp and grasp poses, we move to each pose sequentially by linearly and spherically linearly interpolating to produce a smooth trajectory. An example of a grasp sequence is shown in Figure 1. During grasping, two modules play a critical role in the robustness of LPP: compliant control for grasping (CCG) and tactile regrasping (TR).
Compliant Control for Grasping We use a cartesian impedance controller for manipulation. This choice allows us to set the level of compliance of the manipulator at various stages as needed. Before closing the gripper to grasp, we set the translational and rotational stiffness of the manipulator’s controller to minimal values. The low compliance of the robot arm then naturally allows the gripper to comply with the contact constraints and minimize its applied forces against the environment.
Tactile Regrasping Our method assumes that the grasps that we perform are stable, that is, the object does not move or slip once grasped. To mitigate this issue, we use tactile sensing to preemptively detect if a grasp is stable and perform a regrasp if needed.
In Section IV-A, we experimentally investigate the effects of including these modules through ablation studies.
III-A3 Retrieving
We begin the retrieval phase when a valid stable grasp is detected. An image of the retrieval procedure is displayed in Figure 1. We store the pose of the end effector (EE) as well as the respective timestamp throughout the retrieval process. The manipulator is set to be compliant along the rotational axes of the EE when commanded to move to the pregrasp pose. Doing so allows the manipulator to better comply to any unexpected interactions with the environment. Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{c,k}$ centered around a fixed pose above the scene. That is, given a fixed clearance pose, we add random translations sampled from $\mathcal{N}(0,\,\sigma_{tr}^{2})$. We use $\sigma_{tr}$ = 2.5 $\mathrm{cm}$ during data collection. At the end of the retrieval phase, we have an expert retrieval trajectory of length $M$ from grasp pose to clearance pose via a pregrasp pose:
$\tau^{\text{expert}}_{r}=((\mathbf{T}_{0},t_{0}),...,(\mathbf{T}_{M},t_{M})),$ (1)
where $\mathbf{T}_{0}=\mathbf{T}_{g,k}$, $\mathbf{T}_{M}=\mathbf{T}_{c,k}$ and $t_{0},...,t_{M}$ are the respective timestamps of the trajectory starting with $t_{0}$ = 0. To match the desired control frequency of our policy, we sample states $\Delta t$ apart to generate a more sparse trajectory $\bar{\tau}^{\text{expert}}_{r}$ of length $\bar{M}<M$. Our visual policies operate at 5 $\mathrm{h}\mathrm{z}$, so we set $\Delta t=$ 0.20 $\mathrm{s}$.
III-A4 Placing by reversing
We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory:
$\bar{\tau}^{\text{expert}}_{p}=(\mathbf{T}_{\Delta t_{\bar{M}}},...,\mathbf{T}_{\Delta t_{0}}).$ (2)
An example of a place trajectory is shown in Figure 1. Additionally, we convert the global poses to relative poses in order to use state differences as expert actions:
$\mathbf{T}_{\Delta\text{cmd},i}=\mathbf{T}^{-1}_{\Delta t_{\bar{M}-i}}\mathbf{T}_{\Delta t_{\bar{M}-i-1}},$ (3)
where $i\in\{1,...,\bar{M}\}$ and $\mathbf{T}_{\Delta\text{cmd},i}\in SE(3)$ is a relative change in EE pose. In practice, we use six-dimensional vectors $[\mathbf{t},\boldsymbol{\theta}]^{T}$ to represent poses, where $\mathbf{t}\in\mathbb{R}^{3}$ is a translation vector and $\boldsymbol{\theta}\in\mathbb{R}^{3}$ is a rotation vector. The rotation vector is defined as $\boldsymbol{\theta}=\theta\mathbf{e}$, where $\theta\in\mathbb{R}^{1}$ is the rotation angle and $\mathbf{e}\in\mathbb{R}^{3}$ is the rotation axis. We define the expert action as $\mathbf{a}_{i}=(\mathbf{a}_{\Delta\text{cmd},i},a_{\text{gripper},i})$ where $\mathbf{a}_{\Delta\text{cmd},i}=[\mathbf{t}_{i},\boldsymbol{\theta}_{i}]^{T}$ is the six-dimensional vector representation of the pose $\mathbf{T}_{\Delta\text{cmd},i}$ and $a_{\text{gripper},i}\in\{0,1\}$ is a binary variable describing the gripper command (open or close). To collect an episode of expert data, we perform a rollout of the downsampled place commands, $\{\mathbf{a}_{i}\}_{i=0}^{\bar{M}}$, and store the sequences of observations and actions as an expert trajectory. We use RGB images from the wrist camera to form the robot’s observation space $\mathbf{o}_{i}$, where $\mathbf{o}_{i}\in\mathbb{R}^{w\times h\times 3}$. We downsample the images to a resolution of $128\times 128$. During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the following gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}\}$). Finally, at the end of the trajectory, we command the robot to keep the EE in its current pose and to open the gripper for N time steps (i.e. $\mathbf{a}_{\Delta\text{cmd},j}=[0,0,0,0,0,0]^{T}$ and $a_{\text{gripper},j}=0$ for all $j\in\{\bar{M}+1,...,\bar{M}+N\}$). We use $N=5$, which gives the robot an adequate amount of time to open the gripper and release the object. A single expert demonstration or trajectory is then defined as a set of training tuples:
$\tau^{expert}=\{\mathbf{o}_{t},\mathbf{a}_{t},\mathbf{o}_{t+1}\}_{t=0}^{T},$ (4)
with a total length of $T=\bar{M}+N$.
III-B Noise-Augmented Data Collection
To increase the coverage of our demonstrations and train more robust policies, we perturb the first $75\%$ of the poses with noise. We draw inspiration from previous works that have investigated the use of noise injection for both the policy observations [29, 30], and actions or controls [31]. We call this variant noise-augmented data collection. We perturb the translation of the pose $\mathbf{t}$ as follows:
$\mathbf{t}_{perturbed}=\mathbf{t}+\delta\mathbf{t},$ (5)
where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$. Similarly, we perturb the rotation as follows:
$\boldsymbol{\theta}_{perturbed}=(\theta+\delta\theta)(\mathbf{e}+\delta\mathbf{e}),$ (6)
where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$. We find $\sigma_{t}=0.5$\mathrm{c}\mathrm{m}$$, $\sigma_{e}=0.5$\mathrm{c}\mathrm{m}$$ and $\sigma_{\theta}=$$$ 0.5  ° to be reasonable values. We evaluate the effects of this augmentation on policy performance in Section IV-B
III-C Policy Learning
We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}|\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by LPP. We use behavioral cloning (BC) [32] to train our policy with a likelihood-based loss or objective function. Given a dataset of $N_{train}$ demonstration trajectories, our loss function is then:
$\mathcal{L}=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\sum_{t=0}^{T}-\log{\pi_{\phi}(\mathbf{a}_{t}|\mathbf{o}_{t})}.$ (7)
The main backbone of our policy network consists of a convolutional neural network (CNN) based on the ResNet18 architecture [33], followed by a multilayer perceptron (MLP) layer that maps the CNN output into the parameters of the distribution of actions. In this work, we consider two representations for the action distribution $\pi_{\phi}(\mathbf{a}|\mathbf{o})$: a unimodal Gaussian and a multimodal mixture of Gaussians. In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [34, 35, 36], by setting the standard deviation of the Gaussian distributions to be the arbitrarily small value of $61e-$ 6. Note that the unimodal Gaussian is then equivalent to a standard deterministic policy trained with a mean squared error (MSE) loss. As commonly done [37], we stack a history of RGB frames into sequences of length four for our observations.
IV Experiments
[…]
Figure 3: Sequence of images from roll outs of place policies trained using data collected with LPP. The policies are able to place objects of varying properties in the scene using images from the wrist camera directly.
We use LPP to train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Figure 3. In both cases, the place policy has to account for the weight and shape of each item, the initial grasp of the object and the state of the current scene to decide where and how to place (e.g., place angle). In this section, we present our quantitative results for the dishrack loading task as it better represents the type of task, with contact constraints from the environment, that LPP was designed for. We study the robustness of LPP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B and the quality the policies trained with data from LPP and kinesthetic teaching in Section IV-C.
IV-A LPP Data Collection Robustness
Figure 4: Tactile images from before and after a tactile regrasp (TR). Left: the plate’s edge is seen in the middle of the tactile image indicating a shallow grasp. Right: after a regrasp, the edge of the plate has shifted to a desired spot indicating a stable grasp.
Figure 5: Ablation study on compliant control for grasping (CCG) and tactile regrasp (TR). We measure the average number of episodes that can be collected autonomously based on the number of failures. Both CCG and TR play a crucial role in reducing the number of failures to zero, which allows us to collect the necessary amount of data seamlessly.
We require a robust data collection loop to collect a large amount of place demonstrations autonomously with LPP. We study the effects of compliant control for grasping (CCG) and tactile regrasp (TR), as introduced in Section III, on the robustness of LPP.
We observed two main causes that led to misplaced plates during data collection. First, the manipulator would stiffly grasp the object causing it to move and generate a large amount of force preload against the dishrack. During retrieval, significant object motion occurred at the moment when the object’s contact force with the environment was released. This unexpected shift in the object pose lead to less accurate placements since the reversal procedure no longer resulted in correct alignment between the object and the dishrack slot. With LPP, we mitigate this failure mode with CCG where we decrease the translational and rotational stiffness of the manipulator controller right before grasping the object. The second failure mode involved unstable grasps that lead to large relative motion between the object and the gripper. With LPP, we mitigate unstable grasps using TR to preemptively detect if a grasp will lead to relative object motion and to regrasp if needed. Based on empirical data, we find that unstable grasps typically occurred with shallow grasps. To this end, we design our regrasping module to recognize edges and calculate the relative change in EE pose required to recover a deep enough stable grasp.
[…]
Figure 6: Visualization of edges in the tactile image from grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
To determine the position of an edge, we fit an ellipse to the detected contact region. If an edge is detected, we calculate the difference between the ideal and detected depth of the edge and use this to calculate the commanded change in EE pose. An example of this is shown in Figure 4. While not all graspable objects contain edges, a large subset of items do as shown in Figure 6.
We conduct an ablation study on CCG and TR where we run self-supervised data collection for a total of 128 episodes and record the number of failures. We plot the results in Figure 5. The naive approach without CCG and without TR fails a total of 15 times, which roughly translates to an average of nine episodes collected autonomously before a human intervention is needed. The amount of failures drop to three with the addition of CCG, which raises the average to $43$ episodes collected before each intervention. Finally, with both CCG and TR, we are able to successfully collect 128 episodes without any human intervention.
IV-B Noise Augmentation Ablation
TABLE I: Success rates for a model trained on a dataset collected with and without noise augmentation using two different policy action representations. The noise-augmented data collection procedure improves the performance of both deterministic and Gaussian mixture policies.
Action Representation Trained with Noise Aug. Data Avg. Success Rate ($\uparrow$) Total Successes ($\uparrow$)
Deterministic No 71.67(5.93) 57/80
Yes 81.11(5.50) 56/70
Gaussian Mixture No 72.22(5.67) 50/70
Yes 87.78(3.93) 62/70
[…]
We hypothesize that adding noise further accentuates the multimodal nature of the data (i.e., multiple valid actions are available for roughly the same observation) and that a mixture representation for the action allows the policy to better capture this by not having to average over multiple possible expert actions [38]. We also qualitatively observe that policies trained with noise-augmented datasets tend to be better at re-adjusting the object with respect to the goal.
IV-C Comparison to Kinesthetic Teaching
Figure 7: Comparison of success rates of policies trained using a dataset collected by LPP and a traditional kinesthetic teaching approach. We test models trained on an increasing amount of demonstration data. Policies trained on data collected by LPP outperform those trained on data collected from kinesthetic teaching in almost all dataset sizes. The shaded region consists of one standard deviation.
We compare the performance of policies trained using demonstration data from LPP and from kinesthetic teaching. In particular, we investigate whether the quality of the demonstrations differ. To compare, we train two separate policies, both sharing the architecture outlined in Section III-C, on two separate datasets: one collected with LPP and another collected with kinesthetic teaching [39]. We use a Gaussian mixture model as the action representation of both policies. For LPP, we use the noise-augmented variant since it was the best-performing model, as shown in Section IV-B.
[…]
We hypothesize that the kinesthetic policy performs more poorly than the LPP policy because the overall quality of the human demonstrations is lower than the programmatic or machine-generated demonstrations produced by LPP. Previous work has studied the difficulties and challenges of training policies using human demonstrations when compared to programmatic or machine-generated demonstrations [40, 41, 42]. In particular, human demonstrations can be sub-optimal due to human mistakes (e.g., large variations in trajectory length, unwanted movements, and remaining idle for some time to think).
V Limitations and Future Work
The quality and diversity of the collected grasps in LPP are partly determined by the viewpoint of the camera. In this work, we empirically select a viewpoint that is adequate for our use-case. It would be interesting to allow the manipulator to explore and find vantage points for better grasp generation. Furthermore, while tactile sensing is used to improved the robustness of data collection, it could also be used as an extra modality for the policy.
VI Conclusion
We presented LPP, a self-supervised data collection method to collect expert demonstrations for placement tasks. LPP leverages recent advancement in grasp planners and the symmetry between picking and placing to autonomously collect placing demonstrations. LPP is able to robustly collect a large amount of data consistently without human intervention by making use of compliant manipulator control and regrasping based on tactile sensing. We showed that the quality of the demonstrations produced by LPP surpasses that of human demonstrations collected by kinesthetic teaching. LPP provides a promising and simple direction for real world robotic placing demonstration data collection with minimal human labour.